OpenStudy (anonymous):

Select all of the following expressions that are equivalent to 0.25. 1. (0.5)2 2. (-0.2) + (-0.05) 3. |-3 + 2.75| 4. -(-0.25) 5. 2x - 1.75x

OpenStudy (math&ing001):

1. (0.5)2=1 2. (-0.2) + (-0.05) = -0.25 3. |-3 + 2.75| = |-0.25| = 0.25 4. -(-0.25) = 0.25 5. 2x - 1.75x = 0.25x

OpenStudy (anonymous):

thank you @math&ing001

OpenStudy (math&ing001):

you're welcome